OpenStudy (anonymous):

A chemist mixed some 15% saline solution with some 20% saline solution to obtain 200 mL of an 18% saline solution. How many milliliters of the 15% saline solution did the chemist use in the mixture?

OpenStudy (anonymous):

Use variables x and y. .15x + .2y = 200(.18)

OpenStudy (anonymous):

x + y = 200 .15x + .2y = .18(200) = .36 .15x - .15y =.15(200) = .-30 .05y = 6

OpenStudy (anonymous):

Ohhhhhhhhhh....Thanks!

OpenStudy (anonymous):

.05y = 6 y = 120 x = 200 - y = 200 - 120 x = 80 ml